OpenStudy (rock_mit182):

Find the rational exponent. • Simplify each expression using the rules of exponents and examine the steps you are taking.

OpenStudy (rock_mit182):

\[\frac{ a ^{\frac{ 1 }{ 2 }} }{ a ^{2} }\]

OpenStudy (rock_mit182):

\[\frac{ 1 }{ a ^{-\frac{ 1 }{ 2 }} }\]

OpenStudy (rock_mit182):

\[\frac{ 1 }{ a ^{2} a ^{-\frac{ 1 }{ 2 }}}\]

OpenStudy (rock_mit182):

\[\frac{ 1 }{ a ^{2-\frac{ 1 }{ 2 }} } = \frac{ 1 }{ a ^{\frac{ 3 }{ 2 }} } = 1/\sqrt{a ^{3}}\]

OpenStudy (tkhunny):

"Simplify" is a funny thing. It's particularly funny how "simplify" can mean to make a total mess and call it better. I would recommend \(a^{-3/2}\) on this one.

OpenStudy (mathmate):

Rule of exponent you need on this one is: \(\huge \frac{a^m}{a^n}=a^{m-n}\)

OpenStudy (rock_mit182):

@mathmate which one is the simplest form

OpenStudy (rock_mit182):

a ^-3/2

OpenStudy (rock_mit182):

?

OpenStudy (rock_mit182):

Please help me out. i need to find out the simplest form

OpenStudy (mathmate):

@rock_mit182 Yes, your answer is correct, here's why: \(\huge \frac{a^{1/2}}{a^2}=a^{1/2-2}=a^{-3/2}\)

OpenStudy (rock_mit182):

i don't have to rationalize ?

OpenStudy (mathmate):

Yes, you do have to rationalize, but only when you need to. You do not have a radical in the denominator, just a negative fractional exponent, so no need.

OpenStudy (rock_mit182):

ok thanks

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (rock_mit182):

\[(\frac{ a ^{\frac{ 1 }{ 4 }} }{ a })^{2}\]

OpenStudy (rock_mit182):

\[\frac{ (a ^{\frac{ 1 }{ 4 }})^{2} }{ (a)^{2} }\]

OpenStudy (mathmate):

Use the law of exponents: \(\large (a^m)^n=a^{mn}\)

OpenStudy (tkhunny):

There is no such thing as a "simplest form". It's subjective.